# Ameliorate Spurious Correlations in Dataset Condensation

### v1

@@ Abstract
Dataset Condensation has emerged as a technique for compressing large datasets into smaller synthetic counterparts, facilitating downstream training tasks. In this paper, we study the impact of bias inside the original dataset on the performance of dataset condensation. With a comprehensive empirical evaluation on canonical datasets with color, corruption and background biases, we found that color and background biases in the original dataset will be amplified through the condensation process, resulting in a notable decline in the performance of models trained on the condensed dataset, while corruption bias is suppressed through the condensation process. To reduce bias amplification in dataset condensation, we introduce a simple yet highly effective approach based on a sample reweighting scheme utilizing kernel density estimation. Empirical results on multiple real-world and synthetic datasets demonstrate the effectiveness of the proposed method. Notably, on CMNIST with 5% bias-conflict ratio and IPC 50, our method achieves 91.5% test accuracy compared to 23.8% from vanilla DM, boosting the performance by 67.7%, whereas applying state-of-the-art debiasing method on the same dataset only achieves 53.7% accuracy. Our findings highlight the importance of addressing biases in dataset condensation and provide a promising avenue to address bias amplification in the process.
@@ 1 Introduction
Dataset plays a central role in machine learning model performances. The rapidly growing size of contemporary datasets not only poses challenges to data storage and preprocessing, but also makes it increasingly expensive to train machine learning models and design new methods, such as architectures, hyperparameters, and loss functions. As a result, dataset condensation (also known as dataset distillation (Wang et al., 2018)) emerges as a promising direction for solving this issue. Dataset Condensation aims at compressing the original large-scale dataset into a small subset of information-rich examples, enabling models trained on them to achieve competitive performance compared with training on the whole dataset. The distilled dataset with a significantly reduced size can therefore be used to accelerate model training and reduce data storage. Dataset condensation has been shown to benefit a wide range of machine learning tasks, such as Neural Architecture Search (Wang et al., 2021), Federated Learning (Xiong et al., 2023), Continual Learning (Yang et al., 2023; Gu et al., 2023), Graph Compression (Jin et al., 2021; Feng et al., 2023a), and Multimodality (Wu et al., 2023).
Recent dataset condensation methods mainly focus on performance enhancements on standard datasets (Loo et al., 2023; Zhao et al., 2023; Du et al., 2023; Cui et al., 2023). However, this emphasis often overlooks dataset bias, a critical issue in machine learning. Dataset bias (Tommasi et al., 2017) occurs when collected data unintentionally reflects existing biases, leading to skewed predictions and potential ethical concerns. Despite extensive research on bias detection and mitigation strategies in recent years (Sagawa et al., 2019; Li & Vasconcelos, 2019; Nam et al., 2020; Lee et al., 2021; Hwang et al., 2022), the impact of dataset bias on data condensation remains unexplored. Given that a biased synthetic set can lead to inaccurate or unfair decisions, understanding the role of bias in dataset condensation and developing effective mitigation strategies is crucial.
@@
To answer the first question, we assess existing dataset condensation algorithms across several benchmark datasets. Our findings reveal that the condensation process is significantly influenced by the type of bias, with color and background biases being amplified and noise bias suppressed. For datasets exhibiting amplified biases, we found that even state-of-the-art de-biasing training methods are insufficient in restoring the original performance. This highlights the urgency of developing a bias-mitigating dataset condensation algorithm.
To counter bias in condensed datasets, we propose a simple yet effective debiasing algorithm based on data point reweighting. Leveraging the insight that biased data points cluster in the model’s embedding space, we down-weight samples within such clusters using kernel density estimation. This re-weighting rebalances the significance of biased and unbiased samples, ameliorating biases in the condensation process.
@@
    We provide the first study on the impact of biases in dataset condensation process.
@@
    We propose a simple yet effective re-weighting scheme to mitigate biases in two canonical types of dataset condensation methods.
@@ 2 Related Work
2.1 Dataset Condensation
Dataset Condensation (DC) or Distillation (DD) (Wang et al., 2018; Li et al., 2022) aims to compress a large dataset into a small but informative one to achieve competitive performances compared to the whole dataset. Naturally, it’s framed as a bi-level optimization problem using surrogate objectives such as model gradients (Zhao & Bilen, 2021b; Feng et al., 2023b; Du et al., 2024) or training trajectories (Cazenavette et al., 2022; Du et al., 2023; Cui et al., 2023; Guo et al., 2023). The inner loop of the bi-level optimization usually optimizes a surrogate model using the distilled dataset and the outer loop tries to optimize the resulting distilled dataset itself.
@@ 2.3 De-biasing Methods
Note that, although some prior de-biasing methods involve reweighting (Sagawa et al., 2019; Li & Vasconcelos, 2019; Nam et al., 2020), our approach differs for the following reasons: 1) Previous methods (Li & Vasconcelos, 2019; Nam et al., 2020) integrate the reweighting scheme into the final de-biased model’s training, either as a loss or using an auxiliary model, optimizing it alongside the de-biased model. This process does not apply to dataset condensation methods since a well-trained model is not required in the process (e.g., DM uses randomly initialized models, and MTT only matches part of the training trajectories). 2) Methods like (Sagawa et al., 2019) require explicit bias supervision, which may be challenging or infeasible, while our method has no such requirement.
3 The Impact of Bias in Dataset Condensation
In this section, we conduct comprehensive experiments to answer the following question: how does a biased training set influence the distilled data? Will the bias be amplified or suppressed through the condensation process?
@@ 3 The Impact of Bias in Dataset Condensation
We conduct experiments on three representitive dataset distillation algorithms: DSA (Zhao & Bilen, 2021b), DM (Zhao & Bilen, 2021a) and MTT (Cazenavette et al., 2022) as lots of SOTA methods can be used as an add-on module to these methods (Wang et al., 2022; Kim et al., 2022; Liu et al., 2022; Lee et al., 2022). Additionally, we include a baseline approach without dataset condensation, which is equivalent to the case when $\mathcal{F}$ represents an identity transformation. The visualization of bias impacts in dataset condensation can be found in Figure 2.
@@
4 Ameliorate Bias in Dataset Condensation
@@ 4.1 Bias Amelioration Through Re-weighting
In dataset condensation methods such as DM (Zhao & Bilen, 2021a; Zhao et al., 2023), the objective is to match the embeddings generated by the synthetic dataset ($\mathcal{S}$) with the ones generated by the real images ($\mathcal{T}$). The objective function is formulated as below
@@ 4.3 Distance Computation in KDE
4.4 Apply to Other Dataset Condensation Methods
In addition to DM and its follow up works such as DREAM (Liu et al., 2023) and IDM (Zhao et al., 2023), our method can also be easily applied to other dataset condensation methods. Here we use DSA (Zhao & Bilen, 2021b, 2023; Kim et al., 2022) as an example which synthesizes data by matching the gradients between surrogate models trained with real data and synthetic data. Generally, gradient matching based methods can be formulated as:
@@ 5.2 Experimental Results
We evaluate the performance of 4 representative dataset condensation methods: DM (Zhao & Bilen, 2021a) , DSA (Zhao & Bilen, 2021b), DREAM (Liu et al., 2023) and IDM (Zhao et al., 2023). We then use the proposed algorithm to improve DM and DSA algorithms in the experiments. Note that our algorithm can also be applied to other existing condensation methods, as discussed in Section A.1. Also, see Section A.7 for a qualitative analysis and Section A.2 for specific hyper-parameter settings.
@@ 5.3 Runtime Analysis
Our method has a similar runtime overhead to the state-of-the-art de-biasing method SelecMix (Hwang et al., 2022). It consists of two parts: 1) training a ResNet18 (He et al., 2016a) with a 128-dimensional MLP projection head, typically converging within 30 minutes which is around 5% overhead on DSA which usually requires more than 10 hours to finish with IPC 50 on the hardware, and 2) computing pairwise distances and KDE within each training batch, with a time complexity of $\mathcal{O}(B^{2}d)$, where $B$ is the batch size (usually 256) and $d$ is the projection dimension (128 in our case). With DSA and IPC 50, it takes 55 seconds for 10 iterations without KDE and 65 seconds with KDE, resulting in an 18.2% overhead. Note that, unlike SelecMix which updates the SupCon model during training and repeatedly recomputes pairwise distances, our pretrained SupCon model allows us to cache and reuse pairwise distances throughout the entire condensation process to mitigate this overhead. More details can be found in Section A.12.
@@ 6 Ablation Study
Remark 2 Our experiments reveal that biased features will still dominate even when using de-biased surrogate models during dataset condensation. This may be due to sub-optimal bias feature disentanglement and less stable matching trajectories.
@@
Remark 4 Creating an unbiased dataset first and then applying dataset condensation methods on it also greatly boost the model performance. However it still falls short by a large margin compared to our method.
@@ 7 Conclusion
This paper conducts the first analysis of dataset bias in dataset distillation. Our findings show that bias type greatly influences distillation behavior (amplification vs. suppression). Then we introduce a debiasing method using re-weight and kernel density estimation which substantially reduces retained bias in distilled datasets. We assess our debiasing method on various benchmark datasets with different bias ratios and IPC values and empirically verify the effectiveness of our method. In summary, our study offers insights into bias in dataset condensation, presents a practical algorithm for better performance and paves the way for future research on bias mitigation.
Limitations. Although many dataset condensation methods rely on the matching of real and synthetic dataset through carefully designed objective functions, there are methods such as MTT where our method cannot be easily applied due to the use of expert training trajectories instead of real images directly. This will be the focus of our future research.
@@
Our paper will have the following positive social impacts: 1) We are the first revealing bias amplification effect in data condensation processes, which can prevent the misuse of such technique. 2) We propose a simple yet effective way to mitigate bias when data condensation is used, which can lead to unbiased decisions when employing those models.

### v2

# Mitigating Bias in Dataset Distillation

@@ Abstract
Dataset Distillation has emerged as a technique for compressing large datasets into smaller synthetic counterparts, facilitating downstream training tasks. In this paper, we study the impact of bias inside the original dataset on the performance of dataset distillation. With a comprehensive empirical evaluation on canonical datasets with color, corruption and background biases, we found that color and background biases in the original dataset will be amplified through the distillation process, resulting in a notable decline in the performance of models trained on the distilled dataset, while corruption bias is suppressed through the distillation process. To reduce bias amplification in dataset distillation, we introduce a simple yet highly effective approach based on a sample reweighting scheme utilizing kernel density estimation. Empirical results on multiple real-world and synthetic datasets demonstrate the effectiveness of the proposed method. Notably, on CMNIST with 5% bias-conflict ratio and IPC 50, our method achieves 91.5% test accuracy compared to 23.8% from vanilla DM, boosting the performance by 67.7%, whereas applying state-of-the-art debiasing method on the same dataset only achieves 53.7% accuracy. Our findings highlight the importance of addressing biases in dataset distillation and provide a promising avenue to address bias amplification in the process.
@@ 1 Introduction
Dataset plays a central role in machine learning model performances. The rapidly growing size of contemporary datasets not only poses challenges to data storage and preprocessing, but also makes it increasingly expensive to train machine learning models and design new methods, such as architectures, hyperparameters, and loss functions. As a result, dataset distillation (also known as dataset condensation (Wang et al., 2018)) emerges as a promising direction for solving this issue. Dataset Distillation aims at compressing the original large-scale dataset into a small subset of information-rich examples, enabling models trained on them to achieve competitive performance compared with training on the whole dataset. The distilled dataset with a significantly reduced size can therefore be used to accelerate model training and reduce data storage. Dataset distillation has been shown to benefit a wide range of machine learning tasks, such as Neural Architecture Search (Wang et al., 2021), Federated Learning (Xiong et al., 2023), Continual Learning (Yang et al., 2023; Gu et al., 2023), Graph Compression (Jin et al., 2021; Feng et al., 2023a), and Multimodality (Wu et al., 2023).
Recent dataset distillation methods mainly focus on performance enhancements on standard datasets (Loo et al., 2023; Zhao et al., 2023; Du et al., 2023; Cui et al., 2023). However, this emphasis often overlooks dataset bias, a critical issue in machine learning. Dataset bias (Tommasi et al., 2017) occurs when collected data unintentionally reflects existing biases, leading to skewed predictions and potential ethical concerns. Despite extensive research on bias detection and mitigation strategies in recent years (Sagawa et al., 2019; Li & Vasconcelos, 2019; Nam et al., 2020; Lee et al., 2021; Hwang et al., 2022), the impact of dataset bias on data distillation remains unexplored. Given that a biased synthetic set can lead to inaccurate or unfair decisions, understanding the role of bias in dataset distillation and developing effective mitigation strategies is crucial.
@@
To answer the first question, we assess existing dataset distillation algorithms across several benchmark datasets. Our findings reveal that the distillation process is significantly influenced by the type of bias, with color and background biases being amplified and noise bias suppressed. For datasets exhibiting amplified biases, we found that even state-of-the-art de-biasing training methods are insufficient in restoring the original performance. This highlights the urgency of developing a bias-mitigating dataset distillation algorithm.
To counter bias in distilled datasets, we propose a simple yet effective debiasing algorithm based on data point reweighting. Leveraging the insight that biased data points cluster in the model’s embedding space, we down-weight samples within such clusters using kernel density estimation. This re-weighting rebalances the significance of biased and unbiased samples, ameliorating biases in the distillation process.
@@
    We provide the first study on the impact of biases in dataset distillation process.
@@
    We propose a simple yet effective re-weighting scheme to mitigate biases in two canonical types of dataset distillation methods.
@@ 2 Related Work
2.1 Dataset Distillation
Dataset Distillation (DD) or Condensation (DC) (Wang et al., 2018; Li et al., 2022) aims to compress a large dataset into a small but informative one to achieve competitive performances compared to the whole dataset. Naturally, it’s framed as a bi-level optimization problem using surrogate objectives such as model gradients (Zhao & Bilen, 2021b; Feng et al., 2023b; Du et al., 2024) or training trajectories (Cazenavette et al., 2022; Du et al., 2023; Cui et al., 2023; Guo et al., 2023). The inner loop of the bi-level optimization usually optimizes a surrogate model using the distilled dataset and the outer loop tries to optimize the resulting distilled dataset itself.
@@ 2.3 De-biasing Methods
Note that, although some prior de-biasing methods involve reweighting (Sagawa et al., 2019; Li & Vasconcelos, 2019; Nam et al., 2020), our approach differs for the following reasons: 1) Previous methods (Li & Vasconcelos, 2019; Nam et al., 2020) integrate the reweighting scheme into the final de-biased model’s training, either as a loss or using an auxiliary model, optimizing it alongside the de-biased model. This process does not apply to dataset distillation methods since a well-trained model is not required in the process (e.g., DM uses randomly initialized models, and MTT only matches part of the training trajectories). 2) Methods like (Sagawa et al., 2019) require explicit bias supervision, which may be challenging or infeasible, while our method has no such requirement.
3 The Impact of Bias in Dataset Distillation
In this section, we conduct comprehensive experiments to answer the following question: how does a biased training set influence the distilled data? Will the bias be amplified or suppressed through the distillation process?
@@ 3 The Impact of Bias in Dataset Distillation
We conduct experiments on three representitive dataset distillation algorithms: DSA (Zhao & Bilen, 2021b), DM (Zhao & Bilen, 2021a) and MTT (Cazenavette et al., 2022) as lots of SOTA methods can be used as an add-on module to these methods (Wang et al., 2022; Kim et al., 2022; Liu et al., 2022; Lee et al., 2022). Additionally, we include a baseline approach without dataset distillation, which is equivalent to the case when $\mathcal{F}$ represents an identity transformation. The visualization of bias impacts in dataset distillation can be found in Figure 2.
@@
4 Ameliorate Bias in Dataset Distillation
@@ 4.1 Bias Amelioration Through Re-weighting
In dataset distillation methods such as DM (Zhao & Bilen, 2021a; Zhao et al., 2023), the objective is to match the embeddings generated by the synthetic dataset ($\mathcal{S}$) with the ones generated by the real images ($\mathcal{T}$). The objective function is formulated as below
@@ 4.3 Distance Computation in KDE
4.4 Apply to Other Dataset Distillation Methods
In addition to DM and its follow up works such as DREAM (Liu et al., 2023) and IDM (Zhao et al., 2023), our method can also be easily applied to other dataset distillation methods. Here we use DSA (Zhao & Bilen, 2021b, 2023; Kim et al., 2022) as an example which synthesizes data by matching the gradients between surrogate models trained with real data and synthetic data. Generally, gradient matching based methods can be formulated as:
@@ 5.2 Experimental Results
We evaluate the performance of 4 representative dataset distillation methods: DM (Zhao & Bilen, 2021a) , DSA (Zhao & Bilen, 2021b), DREAM (Liu et al., 2023) and IDM (Zhao et al., 2023). We then use the proposed algorithm to improve DM and DSA algorithms in the experiments. Note that our algorithm can also be applied to other existing distillation methods, as discussed in Section A.1. Also, see Section A.7 for a qualitative analysis and Section A.2 for specific hyper-parameter settings.
@@ 5.3 Runtime Analysis
Our method has a similar runtime overhead to the state-of-the-art de-biasing method SelecMix (Hwang et al., 2022). It consists of two parts: 1) training a ResNet18 (He et al., 2016a) with a 128-dimensional MLP projection head, typically converging within 30 minutes which is around 5% overhead on DSA which usually requires more than 10 hours to finish with IPC 50 on the hardware, and 2) computing pairwise distances and KDE within each training batch, with a time complexity of $\mathcal{O}(B^{2}d)$, where $B$ is the batch size (usually 256) and $d$ is the projection dimension (128 in our case). With DSA and IPC 50, it takes 55 seconds for 10 iterations without KDE and 65 seconds with KDE, resulting in an 18.2% overhead. Note that, unlike SelecMix which updates the SupCon model during training and repeatedly recomputes pairwise distances, our pretrained SupCon model allows us to cache and reuse pairwise distances throughout the entire distillation process to mitigate this overhead. More details can be found in Section A.12.
@@ 6 Ablation Study
Remark 2 Our experiments reveal that biased features will still dominate even when using de-biased surrogate models during dataset distillation. This may be due to sub-optimal bias feature disentanglement and less stable matching trajectories.
@@
Remark 4 Creating an unbiased dataset first and then applying dataset distillation methods on it also greatly boost the model performance. However it still falls short by a large margin compared to our method.
@@ 7 Conclusion
This paper conducts the first analysis of dataset bias in dataset distillation. Our findings show that bias type greatly influences distillation behavior (amplification vs. suppression). Then we introduce a debiasing method using re-weight and kernel density estimation which substantially reduces retained bias in distilled datasets. We assess our debiasing method on various benchmark datasets with different bias ratios and IPC values and empirically verify the effectiveness of our method. In summary, our study offers insights into bias in dataset distillation, presents a practical algorithm for better performance and paves the way for future research on bias mitigation.
Limitations. Although many dataset distillation methods rely on the matching of real and synthetic dataset through carefully designed objective functions, there are methods such as MTT where our method cannot be easily applied due to the use of expert training trajectories instead of real images directly. This will be the focus of our future research.
@@
Our paper will have the following positive social impacts: 1) We are the first revealing bias amplification effect in data distillation processes, which can prevent the misuse of such technique. 2) We propose a simple yet effective way to mitigate bias when data distillation is used, which can lead to unbiased decisions when employing those models.